In 2018, Caterpillar Inc. had about 649 million shares outstanding. Their book value was $24.0 per share, and the market price was $155.30 per share. The company’s balance sheet shows that the company had $23.80 billion of long-term debt, which was currently selling near par value.

a. What was Caterpillar’s book debt-to-value ratio? (Do not round intermediate calculations. Enter your answer as a decimal rounded to 2 decimal places.)

b. What was its market debt-to-value ratio? (Do not round intermediate calculations. Enter your answer as a decimal rounded to 2 decimal places.)

c. Which measure should you use to calculate the company’s cost of capital?
Answer #1

a)

Book value of shares = 649,000,000 * 24 = 15,576,000,000

Total book value = 15,576,000,000 + 23,800,000,000

Total book value = 39,376,000,000

Book debt to value = Debt value / total value

Book debt to value = 23,800,000,000 / 39,376,000,000

Book debt to value = 0.60

b)

Market value of shares = 649,000,000 * 155.3 = 100,789,700,000

Total value = 100,789,700,000 + 23,800,000,000 = 124,589,700,000

Market debt to value = Debt / total value

Market debt to value = 23,800,000,000 /100,789,700,000

Market debt to value = 0.24

c)

Market debt to value should be used.